Hideyuki MICE

Document Sample
Hideyuki MICE Powered By Docstoc
					Analysis for QA
           Hideyuki Sakamoto
             1st October 2007
        MICE Tracker Phone Meeting
• Check the first result of scanning
  – 11 points in one View were scanned.
• Discussion of QA time
• Discussion of considerable problem
                         Hit fiber v.s. Sr90 position
                                                  • Hit fiber is determined by
                                                  crossed events (triplet)
Hit fiber#

                                                  • Peak on hit fiber# is shifting
                                                  one by one as A0 is moving.
                   1 2    3 4   5 6 7 8 9 10 11
                                                  => Scanning is working correctly.

             #99       Position of Sr90 [mm]                                #108

                                         5        6
                   1             2                                  9                10

                   3                 4   7        8                  11
Method to estimate LY

• ADC saturated events are rejected.
• Find out the position of Sr90 by checking “triplet”
• Plot LY on fiber (triplet) for each view.
   – It is better not to require triplet for this plot since,
        • 1. Sample might be decreased by channels with dead or low pe.
        • 2. Events even stopped at first or second view (V,W) are still
          useful for investigating for first views. But for furthest view
          (X), sample is the same as for the triplet if all fiber for each
          view are “alive”.
Location of Sr90 (@A0=-47509)




 ADC distribution under Sr90

V (#99)                        View ordering

X (#109)


X (#110)                           X
Discussion on the fraction of Pedestal

• Events for available (required PE over 3)
   – V #99 22% (48/665)
   – X #109 4% (29/739)
   – W #110 9% (65/718)                 Efficiency is very low….

• Rate of beta-ray(Sr90) might be enough high to hit fiber on
  V view at each trigger.
   – Need to check if fiber# (channel#) under Sr90 is actually triggered.
   – Or, it is due to self-trigger mode?
   – Under checking….. To be reported soon.
    Discussion on #/trigger
11,000 total event
 9,396 events for good readout
     1,535 events for bad timing in good readout
     7,861 events for good timing in good readout
 1,515 events for after rejecting saturated events (accept ADC!=255)
      145 events at one scanning point (A0= -47509mm)
   48 events on V view for after requiring PE over 3 at one scanning point (A0 = -
  166 events for after requiring crossed (triplet) events

    At the current setup, we can only use 48 events for QA for each
    point out of 1k trigger per location.
    => 10k/48 ~ 200 times => 200k triggers per location is needed
We have 2 considerable problems for QA.

  Newest cosmic-ray run (self) in August @IC (RUN41,43-45)

                                                     2. “Second” peak
                                                       on all views


                         10 p.e.
1. Low signal on X view
   •   Same distributions will be expected at all views by MC
   •   This behavior is also produced at Sr90 run.
  Which MCM has second peak?
            LHB                                  RHB

                               ADC                                ADC

    Pedestal & saturated events are not filled

Second peak exists at all of MCM for both boards connected with
Appearing on the specific channel?

              LHB MCM#4


   Second peak exists at all of channels
   How about the previous result?
First cosmic run in May (self)
@IC (RUN6)                          Cosmic(self)@FNAL (RUN91)

          10 p.e.                           10 p.e.
  • Difficult to discuss on LY at X view
  • There was no second peak in IC data!
Discussion on “second peak”
• Temp. Controller. Is the most doubtful.
   – IC => Oxford temp. control
   – FNAL => Temp. controller on RHB
• Are there any other differences for setup?

• Take another cosmic-ray runs.
   – 1. take with current setup again.
       • check if result of newest will be repeated
   – 2. use on-board temp. controller instead of OXFORD controller.
       • check if result (previous one) will be repeated (of course, we will see
         the “shoulder noise”)
• If both of results will be , then we will need to think of
  “heater” setup again.
Discussion on low PE at X view

• Low PE at X view is observed in both Sr90
  run and newest cosmic-ray run.
• Need to check on waveguide/connection
  – Swapping waveguide of internal/external by
    exchanging connectors at side of Station-box.
• Scanning across the view
   – Checking if the scanning is working correctly
   – Check if the low signal on X view is uniform on the plane
       • It was actually confirmed by checking cosmic-ray run.
• Swapping waveguide
   – Check connection loss, etc..
   – External and Internal
• Take cosmic-ray run again
   – Check the repeat
• Take cosmic-ray run without OXFORD temp. control
   – Check the repeat
• Scanning is working correctly.
• Need solution to decrease QA time
  – Investigating the large fraction of pedestal.
• Low signal on X and “second” peak on all
  will be investigated.

xeniawinifred zoe xeniawinifred zoe not http://
About I am optimistic girl.